Session: Oct 2009 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M. Roma at Sun Nov 8 04:18:41 2009) Failure : RECO
For problems to the MK5, I have not recorded scan from 01:00UT to end.

Onsala (Michael Lindqvist at Sun Nov 8 07:40:21 2009) Minor failures :
We have a strange problem with the noise diode. Otherwise no known problems.

Westerbork (Geert Kuper at Sun Nov 8 10:32:12 2009) Minor failures : PHAS
Wb used in phased array mode. For the 2 MHz bandwidth part we used 2.5 MHz filters. (We do not have 2 MHz filters in our backend.) There was a problem phasing up the 1346.49 MHz uppersideband.(LCP+RCP) (our backend subband 3) The other frequencies were ok.

Torun (Kaz Borkowski at Mon Nov 9 09:28:50 2009) Success :
No known problems.

Effelsberg (Uwe Bach at Mon Nov 9 11:38:54 2009) Minor failures :
First two scans on Crab are crap. Just too bright for Eb: overflow in all VCs. Operator had to insert 12 db of attenuation! Attenuation was removed after 1:52 UT for the remaining sources.

Jodrell_Bank (Alastair Gunn at Thu Dec 3 12:01:32 2009) Success :
The Lovell Telescope was used (Jb1). No known problems.